OpenStudy (mtalhahassan2):

A line is perpendicular to the line 2x-4y+7=0 and that passes through the point P(7,2). Determine the equation of this line in Cartesian form.

OpenStudy (baru):

find the slope of the given line by bringing it to the form y=mx+c to find slope of the new line, use this: if m1 and m2 are slopes of two perpendicular lines, then m1*m2=-1

OpenStudy (mtalhahassan2):

so is the direction vector be (4,2) like we are goona flip because its perpendicular

OpenStudy (mtalhahassan2):

@baru @Abhisar @brennerbussy2001

OpenStudy (mtalhahassan2):

@Zale101 @nincompoop @elvisiscool69 @hartnn @They_Call_Me_Narii @lυἶცἶ0210 @Whitemonsterbunny17 @retirEEd

OpenStudy (baru):

if you are doing it that way.. the slope of the new line is \(<f_x,f_y>\)=<2,-4> (because the gradient is perpendicular to the level curve)

OpenStudy (baru):

not slope.. i meant direction vector. direction vector of the new line is <2,-4>

OpenStudy (baru):

by which method are you trying to solve this question? if its by vectors and dot product, you have to flip <2,-4> to get <4,2> because <4,2> is perpendicular to the "required line" then <4,2> \(\cdot\) <x-7 , y-2> = 0 (dot product of perpendicular vectors is zero) expand that to get the required equation

OpenStudy (mtalhahassan2):

wait so the given direction vector is (2,-4) and know they want the line to be perpendicular right ? @baru @YanaSidlinskiy @hartnn @Thadds2003 @Ghostgate @AlexandervonHumboldt2 @nincompoop @Kitten_is_back

OpenStudy (mtalhahassan2):

@Luigi0210

OpenStudy (mtalhahassan2):

@inkyvoyd @ichliebedich @brennerbussy2001 @abi_carter @lυἶცἶ0210

OpenStudy (mtalhahassan2):

@ParthKohli @mukushla

OpenStudy (mtalhahassan2):

@moo963

Parth (parthkohli):

are you in 11th grade?

OpenStudy (mtalhahassan2):

@kayt131

OpenStudy (mtalhahassan2):

@ParthKohli no i am in 12

OpenStudy (anonymous):

hmm uh may i ask, what grade are you in?

OpenStudy (mtalhahassan2):

@kayt131 12

OpenStudy (anonymous):

oh... idk if i can help you sorry, im in 8th but ill still try.

OpenStudy (mtalhahassan2):

@kayt131 its ok

OpenStudy (anonymous):

yeah i don't understand this. I haven't learned this yet... thanks for asking for my help tho, i would have helped if I understood. sorry...

OpenStudy (baru):